### v2

# Good Modelling Software Practices

Abstract
Frequently in socio-environmental sciences, models are used as tools to represent, understand, project and predict the behaviour of these complex systems. Along the modelling chain, Good Modelling Practices have been evolving that ensure—amongst others—that models are transparent and their results replicable. Whenever such models are represented in software, Good Modelling meet Good Software Practices, such as a tractable development workflow, good code, collaborative development and governance, continuous integration and deployment; and they meet Good Scientific Practices, such as attribution of copyrights and acknowledgement of intellectual property, publication of a software paper and archiving. Too often in existing socio-environmental model software, these practices have been regarded as an add-on to be considered at a later stage only; modellers have shied away from publishing their model as open source out of fear that having to add good practices is too demanding. We here argue for making a habit of following a list of simple and not so simple practices early on in the implementation of the model life cycle. We contextualise cherry-picked and hands-on practices for supporting Good Modelling Practice, and we demonstrate their application in the example context of the Viable North Sea fisheries socio-ecological systems model.
keywords:
Good Modelling Practice , Good Software Practice , Good Scientific Practice
journal: Ecological Modelling
[…]
mm[#1]
[…]
1 Introduction
Frequently in socio-environmental sciences, models are used as tools to represent, understand, project and predict the behaviour of these complex systems, and for many more purposes (Epstein 2008; Edmonds et al. 2019). The degree of a model’s formalisation ranges from conceptual to mathematical equations to implementation in software (Romanowska 2015), and–by definition–all of these models are purpose-driven simplifications of the system they represent (Stachowiak 1973). We here concentrate on computational models, i.e. on socio-environmental models implemented in software, and there are many of those out there: Currently the CoMSES Net (2024) lists 1153 agent-based models (ABM); two decades ago, there were 1360 ecological models (Benz, Hoch, and Legović 2001); in 2015, a survey amongst 42 modellers counted 278 aquatic ecosystem models (A. B. G. Janssen et al. 2015).
So computational models are plenty and omnipresent in our field and they have proven their value in being “fruitfully wrong” (Epstein 2008). They often, however, escape strict falsifiability: the code may be verifiable only within certain accuracy ranges; the model may be validated only in site-specific application (Refsgaard and Henriksen 2004). The more useful assessment for its scientificity is that of fitness-for-purpose, given its purpose, scope and assumptions (Edmonds et al. 2019; Hamilton et al. 2022; Wang et al. 2023). Good Modelling Practices (GMP) aim at ensuring this.
Examples of such practices often named are: a clear purpose, a thorough domain understanding, going from simple to complex, ensuring reproducibility, exploring sensitivities and validation with good quality data (e.g., Crout et al. 2008). The first reference to GMP may have been by Smagorinsky (1982), who claimed that “under any circumstance, good modelling practice demands an awareness of the sensitivity [] to parametrization” (p.16). From here on, GMP were elaborated and became widespread in the field of hydrology with its first handbook on the topic (Van Waveren, R et al. 1999); it has since been applied to all areas of socio-environmental sciences and has been adopted in community standards such as the Overview, Design, Detail (ODD) documentation protocol and its derivatives (Grimm et al. 2006, 2010, 2020).
GMP appear as five phases in the model life cycle (Table 1 of Jakeman et al. 2024): (1) Problem scoping, (2) conceptualisation, (3) formulation and evaluation, (4) application, and (5) perpetuation, and the reiteration of these phases. Good Modelling Software Practices (GMSP) are prominent in phases 3–5 and are thus subsumed under GMP. But where GMP is concerned with the reliability and validity of a model—possibly foremost its —GMSP is concerned with the quality and reliability of a model’s software implementation and beyond: a good quality model software is not restricted to good computer code, but will support iterations through the model life cycle and will enable Good Scientific Practice (DFG 2013; All European Academies 2023).
Other scientific and technical fields where software plays a major role developed the concept of Good Software Practice (GSP); it can be traced back to the origins of the Unix system, which has at its core not a monolithic but highly granular structure of many little programs that “do one thing only, and do it well” (Ritchie and Thompson 1974). These little tools should allow to develop new software in a better way, argued B. W. Kernighan and Plauger (1976), for example by following the DRY principle: “Do not repeat yourself”, and to KISS — “keep it simple, stupid!” (Johnson 1960; Raymond 2003).
Cotemporaneously the Free Software movement emerged to emancipate software from the ownership of companies and consider it a public good, to be (1) run for any purpose, (2) studied and modified, (3) distributed, and (4) modified and distributed (Stallman 1983, 1996), and with it the practices to ensure these freedoms in OSS. Educating about GSP became central in projects such as the Software Carpentry (Wilson 2016), highlighting the utility of live coding, pair programming and “open everything”. Plenty of checklists for following GSP have been published (e.g., Artaza et al. 2016; Wilson et al. 2014).
For Open Data, internationally agreed criteria are that they are Findable, Accessible, Interoperable, and Reusable (Wilkinson et al. 2016). Consequently, also the software evaluating these data should be FAIR: (F) easy to find; (A) retrievable via standardised protocols; (I) interoperable with other software by data exchange or an application programming interfaces (API), and (R) understandable, modifiable, and to be built upon, or incorporated into other software (Barker et al. 2022). Research organisations like CoMSES Net educate about FAIR research software, including rich metadata, code management, archiving, documentation, and licensing [^1] https://tobefair.org CoMSES Net initiative to make models FAIR. Fairness should include recognition of the work of the people that develop such modelling software but are often not acknowledged in publications as the major form of scientific output–the research software engineers (RSE, Katz et al. 2019; Hettrick et al. 2022).
Despite many existing GMP and GSP guidelines, however, much of the model source code corpus—roughly 80% (Barton et al. 2022)—is not published at all along with its scientific publication, and for a trivial reason: Barnes (2010)’s survey states that “the code is a little raw” was named as the main reason for not publishing the model. Here we aim to address this fear and help build confidence that the model code is good enough (Wilson 2016; Barnes 2010). Beyond those, we break principles down to concrete tools that implement good practices during the modelling software creation process.
Figure 1: The triangle formed by good practices in modelling, software and research
We start off by motivating each of the good practices and contextualizing them towards the goal of publishing a model software or a scientific result arising from it; we put GMSP at the center of a triangle formed by good practices in modelling, software development, and research (Figure 1). We describe the tools that can be used in a non-exhaustive way covering the entire range of good practices; you may deviate from the tools we selected, or disagree with them, and you may add others or leave some out that we suggested.
2 Good practices for whom?
All models start with a purpose, and the purpose must be stated (Edmonds et al. 2019), that has since long been the number one GMP. It does not hurt to know about your domain, either; speak to other experts, develop a conceptual model, only then start formalizing your model in math and in software (Wang et al. 2023; Romanowska 2015; Grimm et al. 2006) to arrive at your computational model—a purpose-driven and simplified representation of your system in software. This model is by and for yourself, it is for readers and reviewers, and it is by and for your collaborators, constituting three tiers of user groups that are targeted by tools addressing GMSP.
2.1 Single authors – yourself
Socio-environmental modelling software can be created by a single person; in fact, it often is in student or postdoc projects or individual programming sprints (Hettrick et al. 2022). Unfortunately, from a software development perspective, the necessity of having to be a domain expert at the same time means that 90% of scientists are self-taught developers without formal training in software engineering (Wilson 2016). So that one person will not only have to write and read her own code and to apply it for simulations, she will have to align the software development with her research, will have to understand what she did in the past, she will have to retrace the steps that led her to the current state. All of this necessitates to some degree that the work is stored redundantly in backups, and that changes are documented. She would like to ensure that code changes do not break previous work and does that by testing the model after every update (Rosero, Gómez, and Rodríguez 2016).
Authors of scientific models eventually become authors or co-authors of scientific publications arising from or with the help of a model. As scientific authors, they are bound by ethical considerations such as the Guidelines for Safeguarding Good Scientific Practice, among them the unambiguous declaration intellectual property (IP) to the work and ensuring availability of the model for a prolonged period of time (DFG 2013; All European Academies 2023). The IP declaration carries with it the proper acknowledgement of software the model is built on, and respecting the copyrights and permissions pertaining to third-party sources used. The archiving requirement carries with it the obligation to ensure that technical failures or changes in the circumstances of the model author do not lead to the loss of the model: decentralised backups or public repositories help.
2.2 Reviewers and readers
If the model is to be used to produce scientific results subject to peer review, the single person will have to ensure reproducibility of results. She will have to subject it to an editor or a reviewer, thus make it readable and understandable, and document it. And sometimes, she might be asked to support the reviewers in executing the model somewhere outside her own computer infrastructure. When feedback comes, there should be a platform to file the individual concerns and address them.
To be published in a scientific journal, reviewers need to be able to access and understand the model. Under the hood, the write-good-code advice is old and simple: “Write code that minimises the time it would take someone else to understand it—even if that someone else is you” (Flesch 1950). But it also needs to be easy for a peer to execute your model and to understand its inputs and outputs. Various documentation principles can help to ensure this accessibility, including automatically generated application programming interface (API) references, in-code documentation and the generation of metadata through packaging.
But even before a reviewer invests her time in evaluating a model software, much can be done by the author herself. In fact, where a reviewer might concentrate on model validation, the author could ensure model verification (Sargent 2010). Verification answers the questions: Does the code do what the author intends it to do? To ensure this, it has become GSP to employ unit testing and try to reach a good coverage of the test framework over your entire code base (Ellims, Bridges, and Ince 2006), and to automate verification with the source code managment (SCM) service as Continuous Integration (CI, Shahin, Ali Babar, and Zhu 2017).
How does an author deal with the feedback she receives during a friendly or journal-led review? Often, this comes as an itemised list of points to address; as such it is in an ideal form to be converted to “tickets” or “issues” in the SCM service, or a dedicated issue tracker system linked to the model software. Improvements to the model code can then be tied to the issue tracker, transparently documenting the resolution of those issues, and helping to formulate the rebuttal to reviewer critique.
2.3 Collaborators
If at least one other person is using the model, the permissions—also known as the license terms—become pertinent. This other user needs a way to communicate with the developer, for feature requests or for reporting bugs. If that person intents to improve on your work, the permissions become more important, needing contributor agreements and codes of conduct. How are decisions made about the now joint modelling software?—some form of governance model needs to be established. With the growth of a community, even a community management system might be required, with granular access, distributed roles, and fine-grained permissions.
Regardless of the size of the collaboration, structured reviews, pre-commit hooks, and common coding standards can be used to maintain high code quality. Software needs to be sustainable as you aim for establishing a larger user base, use it in other scientific software, but most importantly, use the model to support scientific conclusions: the model code is considered primary data and must be available a decade after a scientific publication relying on it (DFG 2013); but will your model be re-runnable and reproducible for that period of time (Benureau and Rougier 2018)? Hardware and software environments change, as might the developer’s focus of work, all contributing to code rot (Liew 2017) or software collapse (Hinsen 2019). Project funding, short-term contracts, and mobility requirements stimulate frequent staff or job rotations. Releasing a model as Free and Open Source Software (FOSS) may be a way to facilitate sustaining a modeling software across those rotations.
It is often desirable to collaboratively develop the software, i.e. involve another person or persons in improving the software. With this collaboration come legal and governance decisions, as well as technical requirements. The legal once concerns copyrights of the different contributors, and often of the employers (research organisations): Many academic institutions do not have yet clear guidelines on the legal aspects of how to contribute to collaborative software or how to accept contributions by other institutions. These can be established in contributor agreements. Collaboration is also am insurance against accidents: the “truck factor” asks: “How many people can get hit by a truck (or bus), before the project becomes unmaintainable?”; the Open Source Security Foundation (2024) gold standard requires that a project’s truck factor is at least three.
For collaborative development with a smaller group, a governance system known as the benevolent dictator is frequently encountered. A benevolent dictator in software development refers to a leadership style where one individual, often the project’s creator or lead developer, has significant control over decision-making processes within the project. Despite holding considerable authority, this individual typically exercises their power with the best interests of the project and its community in mind, hence the term “benevolent.” This leadership model aims to maintain direction and cohesion within the project while still allowing for contributions and feedback from other team members or contributors (Schneider 2022). In scientific projects, the governance and licensing terms for a collaboration can be formulated as part of a Memorandum of Understanding or consortium contract.
3 Tools for Good Modelling Software Practice
The tools described here can roughly be categorised as version control, source code management system, licensing, documentation, packaging, good code, archiving and maintenance, and publication.
3.1 Version control software
A transparent and reproducible distributed SCM (formerly version control system, VCS) is the basis for good software and has been termed “possibly the biggest advance in software development technology” (Spolsky 2010). The currently dominant SCM software is Git [^2] https://git-scm.com OSS distributed VCS. Note: all Uniform Resource Locators (URL) in this manuscript have been last visited and checked on August 20, 2024, originally invented by Linus Torvalds, the creator of Linux. As source code is text, the SCM tracks changes in lines or parts of lines of text. It can be very well used to manage other kinds of changing texts, such as the text of this manuscript. In fact, this manuscript was started with git init; git add manuscript.md; git commit -m "feat: Created manuscript". Some care should be taken with the latter commit messages, as they should be short and descriptive to humans, and at best also machine-interpretable, e.g., by following the conventional commits [^3] https://www.conventionalcommits.org How to add human and machine readable commit meaning recommendations. Learning Git could be considered a valuable investment in yourself for any creative work.
With graphical interfaces to Git, such as Sourcetree [^4] https://www.sourcetreeapp.com Free Git client, or various integrations in text editors, such as Visual Studio Code [^5] https://code.visualstudio.com Cross-platform editor with built-in Git, it is now easy to visually follow the step-wise development and provenance of code (or text documents), go back to points in time or critical development stages, to open experimental deviations from main development (git branch) and combine diverging developments (git merge). Did you mess up? Simply retrace your step back git revert; it helps you even to find in the recorded history those developments where things might have unnoticingly gone wrong with git bisect.
Git and others are most powerful as distributed VCS, in combination with other locations on your own computer, an intranet or the internet, for saving your work in different places, the repositories, while keeping all those versions synchronised. The interaction of two repositories is managed by the unidirectional synchronisations git pull and git push. These commands can be used to synchronise the managed code also across different SCM services, effectively allowing redundant and distributed backups and thus minimizing the risk of losing the software from technical or human errors or the risk of vendor lock-in (Nyman and Lindman 2013).
You must have version control and you must have distributed redundancy.
3.2 Source code management service
The most prominent online SCM service is GitHub [^6] https://github.com Public GitHub, but many academic institutions also offer on-premise or cross-institutional dedicated SCM services, such as the community GitLab of the German Helmholtz Association [^7] https://codebase.helmholtz.cloud.de Community Gitlab of the German Helmholtz Association, for their students and researchers. A good reason to choose GitHub is the higher number of potential contributors on this platform, estimated at roughly 100 million. Many automated tools for supporting software development only work on GitHub. On the other hand, on-premise GitLabs may be preferred by academic institutions, because the code is then hosted in the research centre or by a dedicated subcontracted partner and may offer better data and IP protection; it may also offer more elaborated services or dedicated computing resources.
An SCM service is the entry point for collaborators to contribute, provides a ticketing system and release management, and it offers functionalities for CI and continuous deployment (CD, also known as continuous delivery) of the software. Some SCM services also facilitate project management workflows, such as milestone progress tracking; task boards can be used for task tracking, work assignment or Agile development (Beck et al. 2001).
3.2.1 Ticketing system
Often things do not work right away, or an error is detected in the software. For this, SCM services offer ticketing (also called bug tracker or issue tracking) systems, where one records the occurrence of an error, a bug report, or a wish for future development, a feature request. This works well for a single person, but even better when collaborators and reviewers of the software record their observations on faulty or missing issues with the software on this ticketing system. Git commit messages can be linked to these issues mentioning the hash-prefixed issue number, and they should carry a “fix:” or “feat:” prefix when they resolve a bug or implement a new feature [^8] https://www.conventionalcommits.org How to add human and machine readable commit meaning. Beyond the ticketing system, the SCM service may also offer communication facilities like discussion forums, wikis, mailing list or service desks, which often provide cross-referencing functionality to Git commits and issues.
3.2.2 Continuous integration and deployment
CI is a development practice where developers integrate code into a shared repository frequently, ideally several times a day, but at least on every git push to a repository. Each integration is then verified by an automated build and automated tests to detect errors as quickly as possible—and to correct or recover from them (Shore 2004); the frequent and automated checks reduce the risk of accumulating errors. SCM services like GitLab or GitHub provide such automated integrations, but there are many CI tools available outside the comprehensive SCM services, such as Circle CI [^9] https://circleci.com Dedicated CI utitility. CD is often triggered after the CI ends with success. In this automation, the products of a modelling software can be provisioned, such as a complete binary package for download, an updated and nicely formatted documentation, or a suite of simulation results and their statistical evaluation, for example. CD often interacts with other external services to update web pages [^10] https://about.readthedocs.com Popular documentation service ReadTheDocs, to upload to package repositories, or to submit to an archiving service.
3.2.3 Pull requests
Pull requests (PR) offer an SCM service managed way to accept other people’s changes to your software into your development. Collaborators typically duplicate your software in a fork, apply changes locally and then file a PR providing a detailed explanation of the modifications and their purpose. The SCM service allows you to review the changes, possibly ask for further explanations or modifications, to have it automatically tested with CI, and finally to git merge the collaborator’s work, and even to automate acknowledging the contribution. Similarly, a branch-based approach to PRs uses separate feature git branch branches and often allows to create a Changelog [^11] https://keepachangelog.com Keep a ChangeLog principles based on the merging of the feature branch into main.
You must use a collaboration platform, you should use CI, you may use CD.
3.3 Licensing
Model software development is a creative process. It thus constitutes IP and the right to determine how the model software can be used, re-used and shared and modified, i.e. the copyrights. The exact terms are laid down in what is called a license. Without a license there is no permission, so every model software needs a license, and with it the name of the person or organisation holding the copyrights. While some model software may be published under proprietary licenses and without disclosing the source code, the majority of current modelling software is distributed as open source software (OSS [^12] https://opensource.org/osd The Open Source Definition by the Open Source Initiative 2024), and under a permissive or copyleft license, among them the widely used BSD, MIT, Apache and GPL licenses.
There are strategic decisions involved in choosing for copyleft versus permissive licenses, also related to the community in your field and dependent on third-party software used in your modelling software paper. There are tools to support choosing a license [^13] https://choosealicense.com Choose an OSS license, to manage licenses towards better reuse [^14] https://reuse.software Verification for OSS licensing, and to assess the compatibility of different licenses with a project [^15] https://www.fossology.org Checks for OSS license compliance, [^16] https://github.com/oss-review-toolkit/ort OSS Review Toolkit for managing (license) dependencies. Some research software are dual-licensed, to provide at the same time an OSS license for the research community and the public, and a proprietary one for commercial use.
3.3.1 Contributions
With collaboration also comes the obligation to sort out the copyrights evolving from different contributors, who are all creators and thus natural copyright holders (or their organisation). Your contributors may choose to assign their copyrights to you in what is usually called a copyright transfer agreement (CTA) and is well known from the publication process for scientific papers before the Open Access (OA) movement. Alternatively, your contributors may permit you to exercise copyrights arising from their contribution in a separate agreement, a Contributor License Agreement (CLA) or a Fiduciary License Agreement. Project Harmony [^17] https://www.harmonyagreements.org Project Harmony contributor agreements or the Contributor Agreements [^18] https://contributoragreements.org Contributor Agreements infrastructure support the drafting of such agreements.
3.3.2 Collaboration
Collaborative software engineering is an intensely people-oriented activity (Singer, Sim, and Lethbridge 2008); to keep both the software as well as the collaboration healthy, many projects adhere to the Contributor Covenant [^19] https://www.contributor-covenant.org Contributor Covenant codex. This provides guidelines for respectful and constructive behaviour, it prohibits harassment and discrimination, and overall help maintain a positive and welcoming environment.
Traditionally in science, mostly the authors of scientific publications are acknowledged and cited. But for scientific software, there are more roles that do not qualify for authorship: foremost those of the research software engineers, but also, e.g., data managers, administrators, friendly reviewers, HPC support staff. The All Contributors [^20] https://allcontributors.org Contributor acknowledgement automation bot integrates with the GitHub SCM service and facilitates proper acknowledgement of those and other non-author groups.
You must have a license for yourself and contributions and you should have guidelines for collaboration.
3.4 Versioning and Releasing
In contrast to standard journal publications, software is always a work-in-progress (WIP). Even if there are no new features implemented, software needs to receive regular updates because of the rapid technological development – both in hardware as well as software. Without updates to the software, scientific analysis may become irreproducible because the software cannot be used (Liew 2017; Hinsen 2019); but also, heterogeneous and scattered incremental updates of the software may lead to “code decay”, where the global code quality declines despite local improvements (Eick et al. 2001).
This need for continuous maintenance requires tracking the state of software and recording changes to software beyond the capabilities offered by git log. In an SCM, each state can be marked with an identifier git tag that carries a human readable short description as well as version information following a consistent strategy, such as semantic [^21] https://semver.org Semantic versioning or calendar versioning [^22] https://calver.org Calendar versioning. Releases are an elaborated version of tags, and can include further resources, such as additional documentation or pre-compiled binaries. They also integrate with archiving services such as Zenodo upon a new release.
While changes to the source code are tracked in the SCM, the reasoning behind those and the user-focused communication of these changes should be kept in a change log [^23] https://keepachangelog.com Keep a ChangeLog principles, a technology since long enforced by the GNU coding standard (Chen et al. 2004).
You should have versioning.
3.5 Bundling your application
Much of research software, and especially those (like climate models) with a long history, are distributed by giving access to the repository hosted on an online SCM service, or to an archive file that contains the entire source code. Users are then expected to install the (often numerous) requirements, and eventually compile the code and run the model. To improve usability, more state-of-the art software engineering techniques like packaging and containerisation are available (Vanegas Ferro et al. 2022). Not only do they standardise and simplify the installation of the code, but they also improve its findability via machine- and human-readable metadata, support reusability via versioning and ensure proper citeability. We refer to this as bundling your application. This bundling can be integrated in the CD by deploying the bundle in the associated registries of the online SCM services [^24] https://docs.github.com/en/packages, [^25] https://docs.gitlab.com/ee/user/packages/. You can distribute a package of your model software or your model software including a runtime environment
3.5.1 Packages
Packages are commonly used in programming languages to standardise and simplify the installation of software, and to make the software findable via machine- and human-readable metadata. Packages are files that contain other files, most importantly a manifest that tells the package name and version and lists its content. There are language-specific package managers, e.g., for Python, Julia, R, NPM or Fortran, and language-independent package managers, such as Debian’s dpkg or Continuum’s conda [^26] https://conda-forge.org Conda Forge community packages; the latter, however, often depend on the operating system and computer architecture. Packages provide a detailed declaration of dependencies, including version constraints. Tools like versioneer [^27] https://github.com/python-versioneer/python-versioneer Versioning automation provide the possibility to combine the package version with Git tags. Even the smallest analysis scripts can be distributed in form of a package, while some large models can be hard or impossible to package.
Packages can be distributed via package registries to increase visibility and availability of the software (Allen 2019). The metadata contained in the package is consumed by the registries and made available as a catalogue, such as the Python Package Index [^28] https://pypi.org Prominent python package registry.
3.5.2 Container images
Installation can further be simplified for the user by distributing the software as a container image. Like packages, container images are distributed as a single file. They contain an entire operating system layer where the necessary software and its dependencies have been installed already. These images can then be published in container registries to make them reusable. A prominent tool to build containers is Docker [^29] https://www.docker.com/ Docker tool to build containers, a prominent registry the Docker Hub [^30] https://hub.docker.com/ Free to use container image registry, or the built-in container registries of the online SCM services GitHub and GitLab [^31] https://docs.github.com/en/packages, [^32] https://docs.gitlab.com/ee/user/packages/. Apptainer [^33] https://apptainer.org HPC container builder (formerly Singularity) provides container imaging for for High Performance Computing (HPC).
Distributing a model as a container makes it independent of the user’s infrastructure and thus more easily reusable and reproducible (Vanegas Ferro et al. 2022).
You may provide packages or containers.
3.6 Documentation
Proper documentation often determines a model software’s impact and usability. Most developers use inline comments for code, and they are indeed important, but their impact on helping other scientists in contributing or using the software is limited. Above all, proper Readme text files (often in Markdown format as ReadMe.md) are important, even if the author is its sole user. The Readme is usually the first file added to a new project; it is also the first entry point for users of the software to understand what he or she is looking at, as SCM services render it prominently on your project’s start page as a hypertext (HTML) representation. A Readme provides an overview of the modeling software: its purpose, assumptions and scope in narrative form. A well-crafted Readme enhances user experience by providing clear, concise, and relevant information on using the model. It can also include a brief description of the software’s architecture and its main components, as well as installation instructions. It should contain a section about how you want your software to be cited, or where to find the citation instructions (Allen 2019). A helpful template covering these sections is available at makeareadme [^34] https://www.makeareadme.com ReadMe.md template.
A contributing guide fosters a collaborative environment and encourages contributions from other researchers. It often appears as a Contributing.md file or as a section in the Readme; it provides guidelines on how to contribute to the software, the coding standards to follow, and the process for submitting changes. This ensures that the software continues to evolve and improve, benefitting the entire research community.
Tools like Sphinx [^35] https://www.sphinx-doc.org Sphinx documentation tool or MkDocs [^36] https://www.mkdocs.org YAML-based documentation tool automate the process of building documentation and can contribute to creating user-friendly model software more with more ease: By including inline code comments, they support an up-to-date API documentation; additionally, they support doctests, which ensures that the examples in the documentation work as expected, enhancing the reliability of the documentation.
Important sections in any documentation are:
  1. 1.
    Detailed and clear installation instructions, that eliminate guesswork, making the software accessible to a broader audience. These instructions should cover various operating systems and potential issues that might arise during the installation process. They should also specify the prerequisites, such as required libraries or dependencies.
  2. 2.
    A user manual (in its minimal form a getting started guide) is essential for enabling other modellers to rerun and use your code. It provides step-by-step guidance on how to use the software, explains the functionality of different modules, and shows examples of their use. This allows users to evaluate the model’s fitness for purpose and apply it effectively.
  3. 3.
    API documentation provides a detailed description of how the software’s functions work, the parameters and data they accept, and the output they return. This is crucial for users who want to integrate the software into their own code or use it for more complex tasks—or reuse parts of your model in their modelling software.
  4. 4.
    If you aim for contributions by other researchers, the developer manual is a must-have for the onboarding. It contains more detailed information about the framework that you are developing, that do not have place in the contributing guide or user manual.
3.6.1 Self-checks and badges
Part of the documentation may also be devoted to promotion, self-checking, and community building. For these, software badges have become widespread. They are little visual indicators put atop your Readme that inform readers and yourself at first glance about diverse aspects of your model and modelling process (Lee 2018). Things to show are the activity of your development, the status of passing the CI, the percentage of code covered by tests; information about portability and software security, the status of self-assessments, or the publication status and DOI, amongst many others. Badge awarding can be formally categorised and reviewed (NISO 2021), but they can also be created by yourself [^37] https:/shields.io Badge creation; motivating yourself to keep to good practices may be the most important factor in badge awarding, and showing badges has been shown to positively correlate with data sharing (Kidwell et al. 2016). Particularly devoted to an entire array of Good Practices is the Open Source Security Foundation’s OpenSSF badge program [^38] https://www.bestpractices.dev OpenSSF badges. It consists of many self-assessment questions, and at the end reports the practices in your project at a basic, silver, or gold level. Of course, you should show off this badge in your project’s Readme!
You must have documentation in a Readme, you should elaborate documentation.
3.7 Clean and correct code
Software code must be readable fast not only by a computer but also by humans (Flesch 1950). Code and community conventions are key here; e.g., the PEP8 conventions for Python ensure consistency across different softwares and facilitate automated automated code formatting and linting.
3.7.1 Code formatting and linting
Automated formatters help adhere to conventions. Tools like Black [^39] https://black.readthedocs.io Python formatter or isort [^40] https://pycqa.github.io/isort/ Python import sorter adjust (Python) code to meet specific formatting guidelines, eliminating the need for manual formatting and ensuring consistency across the codebase. A general-purpose tool for many file formats and languages is Prettier [^41] https://prettier.io General purpose code prettifier.
Language-specific linters, such as ESLint [^42] https://eslint.org Javascript linter, flake8 or linter-formatters such as Ruff [^43] https://docs.astral.sh/ruff Python linter and formatter, go a step further by running also checks against coding quality rules [^44] https://www.flake8rules.com List of Python coding standards for automated verification. They all provide feedback that can help developers improve their software quality and adhere to best practices.
Linters and formatters can be combined in pre-commit hooks [^45] https://pre-commit.com Triggers checks before git commit. Pre-commit hooks facilitate the formatting and linting process by automatically running a list of tools before each git commit. This ensures that all committed code adheres to the defined conventions and standards.
3.7.2 Code Structure
Good code is probably the oldest of all good practices, going back to the single-purpose Unix tools (Brian W. Kernighan and Pike 1999). This philosophy has found its way into larger programs, which are broken down into manageable and readable building blocks, often functions or modules, that should have a clear and documented way to interact with data and the rest of the code, the API. Even if your code is not a library for use by others, an internally consistent API can help you to encapsulate functionalities, to allow data manipulation only via dedicated accessors, to be explicit about acceptable inputs, and to finally write safer and more correct code.
Separate concerns in the code structure: do not mix functionality with visual appearance, avoid mixing input and output with processing. Try to implement error handling, then try to break your own code. Often, there is no need to code a functionality yourself, better re-use someone else’s working solution from an existing library or piece of code, “be ruthless about eliminating duplication” (Wilson et al. 2017): vice versa, all parts of your modelling software that have an API may also be more easily reusable by others. Together with human-readable and consistent naming practice, which may adhere to community conventions, following this advice should provide simplicity, generality, and clarity, the “bedrock[s] of good software” (Brian W. Kernighan and Pike 1999).
3.7.3 Code verification
As code gets more complex, it is important to be able to ensure that the smaller units of the code function as expected. For this, the concept of unit tests has been developed, where in principle every code unit has a mirrored part either within the code or separate from it in a testing framework. So, every time you write a new function, think about (and implement) how its functionality can be tested. How much of a code is covered with test is called the coverage, and the coverage can help you identify parts of the code that are not well tested. But even if all units of a code are correct, their interaction may not be, and a model might be dysfunctional or produce unreasonable results. To catch these, regression and reproducibility tests (RT) should be regularly run (Rosero, Gómez, and Rodríguez 2016).
You must provide readable code and you should use automated tools for formatting, linting, and verification.
3.8 Archiving
Archiving the source code of the model and post-processing routines is essential Good Scientific Practice (DFG 2013). The simplest solution for open-source projects is using the Software Heritage Project [^46] https://www.softwareheritage.org Software heritage archive crawler. Many SCM services are already crawled by this project, so when you upload your public repository, it will be archived automatically; you can also add your project manually.
A more citeable and permanent archiving solution is guaranteed by uploading the source code and generating a Digital Object Identifier (DOI). There are various institutional platforms that offer this service, the most prominent one is likely Zenodo [^47] https://zenodo.org Zenodo archive. Most of these platforms additionally offer to create versioned DOIs, where each upload gets its own DOI, and there is one single DOI that represents all versions of the model software. Used in combination with releases, each release gets its own DOI and the contributions to the release are citeable separately. Helpers exist to automate this upload, such as Zenodo’s GitHub integration [^48] https://docs.github.com/en/repositories/archiving-a-github-repository/referencing-and-citing-content Zenodo integration, or the hermes workflow [^49] https://github.com/hermes-hmc/hermes Hermes workflow (Druskat et al. 2022).
You must archive and you should do this on a community platform.
3.9 Publish your model
Many scientific journals nowadays request to publish the source code that has been used to generate figures alongside the journal publication. This increases usability and tractability, it does not, however, acknowledge the amount of work that has been put into the software or its documentation (Hettrick 2024); nor does it take into account that the software might be developed further after the paper has been published.
Consequently, software may be published separate from the scientific publication in a dedicated software journal, often allowing to track the development across multiple versions. Such software journals, e.g. the Journal of Open Research Software or the Journal of OSS (JOSS, Smith et al. 2018) are meant to be developer-friendly, i.e. they integrate with the development workflow and tools used to follow GSP.
Such software journal publications are not domain-specific, so they do not generally improve the findability of your software. Therefore, the most important aspect on publishing your code (or its metadata) is to list it in a platform that your colleagues know, such as the communities on Zenodo [^50] https://zenodo.org/communities Zenodo communities. An alternative might be the Research Software Directory [^51] https://research-software-directory.org Helmholtz Research Software directory (Spaaks et al. 2020) that also allows to group software into projects and interests; or, the community targeting ABMs on the CoMSES Net, where code, metadata and documentation is optionally reviewed (M. A. Janssen et al. 2008). An overview of software registries is provided by SciCodes [^52] https://scicodes.net/participants.
You must publish your model software.
3.10 Software development templates
The multitude and the complexity of GSP techniques and tools put a burden on the modeller. Not only does it require acquiring new knowledge, the rapid technological development makes it also difficult to stay up to date. The diverse skill sets of researchers and time constraints further complicates the situation (Wilson 2016).
To jumpstart new modelling software projects GMSP, they can be derived from templates that already provision these techniques from the start of a project (Pirogov 2024). Cruft [^53] https://cruft.github.io/cruft Cruft template manager and cookietemple [^54] https://github.com/cookiejar/cookietemple Template creation keep things up to date behind the scenes. Sommer, Saß, and Benninghoff (2024), e.g., provide a fork- and cruft-based methodology for modelling ecosystems: they provide a standardised setup for post-processing routines, plugins, etc., implementing many of the techniques and tools mentioned in this manuscript.
You may use templates.
4 Good enough modelling software practice - a use case
Viable North Sea (ViNoS) is a socio-ecological model of the German North Sea coastal fisheries (Lemmen et al. 2023, 2024) coded in NetLogo (Wilensky 1999) and embedded in a larger software system containing data, and Python/R data pre- and postprocessing scripts. Its NetLogo source code is broken down into a central NetLogo file and several included object-specific module files. The code, data and scripts are managed by Git, with a primary SCM service on an academic Gitlab [^55] https://codebase.helmholtz.cloud/mussel/netlogo-northsea-species and a secondary one on public Github. On the primary SCM, GitLab issues provide the ticketing system and milestone planning. Upon each git push to the repository, a CI rule triggers license checking, builds Docker images for different versions of the NetLogo IDE, performs unit and replicability testing within these containers. Using them for subsequent CD, a small production simulation generates model results and pushes them to a static web page on the SCM service; the documents provided with the model (ODD, JOSS, this manuscript, and others) are compiled from their version-controlled Markdown sources to pdfs (using pandoc and LaTeX) and uploaded. The CD is integrated with Mkdocs and pushes to a public readthedocs instance for the user guide. On the secondary CI, a release management hooks into Zenodo to provide permanent and DOI citable archives for each model release. Locally, a pre-commit workflow, triggered upon each git commit ensures that the copyrights and license catalogues are complete for each file, that all structured documents comply with their respective type definition, and that python codes conform to PEP8 coding standards. The tools used for this are reuse, black, flake8, more general formatters to remove empty line endings, syntax checkers for metadata in yaml, toml, and json structured formats, and the general-purpose code beautifier prettier.
In the program root folder, a CITATION.cff suggests how to cite the software, a codemeta.json provides package meta information, a ChangeLog.md the user-oriented change information. Every directory contains at least a ReadMe.md. The project implements the Contributor Covenant, its Contributing.md file explains that the best way to contribute is via issues on the SCM service and uses the Project Harmony template for providing the legal infrastructure for the contributor agreements. Badges shown are for CoMSES Net Open Code, Zenodo and JOSS DOIs, an active repository status, the OpenSSF best practices and Contributor Covenant adherence, reuse compliance, prettier code and code quality A.
For reproducibility, a conda environment-dev.yaml for satisfying developer dependencies, as well as a pyproject.toml for the python dependencies, are provided. The model has been archived on Zenodo, and on CoMSES.net together with its ODD (Lemmen et al. 2023). A software paper has been published in JOSS (Lemmen et al. 2024) with a transparent review process [^56] https://github.com/openjournals/joss-reviews/issues/5731; the model has been catalogued in the ABM specific CoMSES.net as well as the Helmholtz research software directory.
5 Conclusion
Good Practices have been maturing in the areas of software, modelling, and research. Adopting all of them for socio-environmental modelling establishes Good Modelling Software Practices: Reproducibility and citeability meet automated verification, rights management and collaboration meet purpose-driven simplification and validation. For this, you must have version control and distributed redundancy, use a collaboration platform, have a license and documentation, provide readable code, and archive and publish your model software. You should use CI, have guidelines for collaboration, elaborate on documentation, employ versioning and releases, use automated tools for formatting, linting, and verification and target a community platform. You may use CD and packages (Figure 2).
Figure 2: Must haves, should haves and may haves in Good Modelling Software Practise
For many scientific model authors, getting stuff done may be more important than documenting it thoroughly, usefulness is more valued than red tape, spontaneous ideas are implemented preferably over those layed down in a management plan; individual agency supersedes organisational processes. Scientific model software development reflects the ideas formulated in agile development: “Individuals and interactions over processes and tools. Working software over comprehensive documentation. [] Responding to change over following a plan. And while the items on the left are valued more, the items on the right are still important” (Beck et al. 2001). In this spirit, Good Modelling Software Practice means that you do not have to do it all at once. The supporting tools are numerous, and most follow a philosophy to do one thing only and to do it well: they can be learned and employed one by one. Publish your modelling software—it is good enough (Allen 2019; Barnes 2010)! By learning and applying one tool at a time, everyone can acquire a corpus of practices that eventually lead to better modelling software—and better research (Katz et al. 2019).
The recent surge in Artificial Intelligence (AI) applications has created an array of emerging good practices in the areas of security (Polemi and Praça 2023), data treatment (Makarov et al. 2021), and in supporting software development (Pudari and Ernst 2023), amongst others. They will need to be negotiated by society but eventually also find their place in supporting Good Modelling Software Practices.